OpenStudy (anonymous):

The hypotenuse of a right triangle is 25 cm, and the shorter leg is 15 cm. Find the length of the other leg.

OpenStudy (anonymous):

\[2\sqrt{5}\] \[\sqrt{10}\] 20cm

OpenStudy (lgbasallote):

pythagorean theorem \[a = \sqrt{c^2 - b^2}\] \[a = \sqrt{25^2 - 15^2}\]

OpenStudy (lgbasallote):

do you know how to do the rest?

OpenStudy (anonymous):

i think so maybe

OpenStudy (lgbasallote):

just call me when you get stuck

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i end up with 20

OpenStudy (lgbasallote):

there you go! :D

OpenStudy (anonymous):

that was easy!! thank you